OpenStudy (anonymous):

Solve the equation for 0 to 360 degrees. cot^2 (x)-csc(x)=1

OpenStudy (mertsj):

\[\frac{\cos ^2x}{\sin ^2x}-\frac{1}{\sin x}=1\] \[\frac {\cos ^2x-\sin x}{\sin ^2x}=1\] \[\frac{1-\sin ^2x-\sin x}{\sin ^2x}=1\] \[1-\sin ^2x-\sin x=\sin ^2x\] \[-2\sin ^2x-\sin x+1=0\] \[\sin ^2x+\sin x-1=0\] \[(2\sin x-1)(\sin x+1)=0\] \[\sin x=\frac{1}{2} or \sin x=-1\] \[x=30^{o}, 150^{o}, 270^{o}\]